# Laura Barros        Aiko Schmeißer

### v1

@@ 5.4 Party disaggregations
Job loss may not only affect overall political participation but also lead to changes in individuals’ support of different political parties. We, therefore, study heterogeneities in the effects on memberships and candidacies by party characteristics, focusing on parties’ ideology, local incumbency, and alignment with labor unions. Estimates from the static DiD model (2) are presented in Table 2, while dynamic event-study results from model (1) are shown in Appendix Figures B.4 - B.6. We also report results from model (2) separately for each of Brazil’s major parties in Appendix Figure B.7.
@@
Ideology. Layoffs may lead individuals to align with parties whose political programs better address worker’s new economic interests, e.g. through redistributive policies. Moreover, even in the absence of economic platforms, extreme parties may be more successful in exploiting individuals’ resentment of the establishment following economic downturns (e.g., Algan et al., 2017; Autor et al., 2020). To analyze whether layoffs lead to ideological shifts, we distinguish left, center, and right parties, following the classification of Colonnelli et al. (2022). [^31] The ideological classification of parties is shown in Appendix Table B.3. Our results, shown in columns (2) to (4) of Table 2, indicate that job loss leads to an increase in party memberships and candidacies across the ideological spectrum. Relative to the control mean, we find an increase of 8.4%, 11.0% and 17.6% in the likelihood of individuals newly affiliating with left, center and right parties, respectively. The effects on candidacy reveal a similar pattern, with likelihood increases of 15.1%, 17.3%, and 21.3% of running for office in a left, center, or right-wing party. Although we observe larger effects for right-wing parties, our results show that center and left-wing parties also receive more members and candidates following layoffs. Interestingly, the event-study results presented in Appendix Figure B.4 show that the higher increase in affiliations with right-wing parties is driven only by the 2020 cycle, the first local election after Bolsonaro took power in 2019. The rise in right-wing support following economic hardships, which is well documented in the literature (e.g., Barros and Santos Silva, 2019; Autor et al., 2020; Dippel et al., 2022), appears to be contingent on the current political climate.
@@
Local incumbency. Next, we examine whether individuals punish the party of the local incumbent government after being laid off. Retrospective voting has been widely documented to rationalize the behavior of voters following economic shocks. In particular, individuals may blame incumbent politicians for local economic policies that result in mass layoffs, or for insufficient support policies that mitigate economic hardship after a layoff (e.g., Tilley et al., 2018; Ahlquist et al., 2020). In columns (5) and (6) of Table 2, we compare the effects for the party of the elected mayor in each individual’s municipality before the layoff in 2008 and all other non-incumbent parties. The results show that job loss leads to increased participation in both incumbent and non-incumbent parties. Relative to the control mean, the effects are considerably stronger for incumbent parties, with an increase of 24.0% in the probability of being affiliated with the party of the incumbent and an increase of 43.7% in the likelihood of running for office in the party of the incumbent (as opposed to 11.6% and 16.7% in non-incumbent parties). The results do not support incumbent punishment as a main driver of our results.
@@
Union affiliation. Lastly, we investigate whether there are heterogeneities depending on parties’ alignment with labor unions. Unions can influence employees’ political behavior through various activities, such as educating, informing, and mobilizing their members, endorsing selected candidates and policies, and organizing get-out-the-vote campaigns (e.g., Kim and Margalit, 2017; Matzat and Schmeißer, 2023). In Brazil, labor unions have historically been linked to left and centrist parties and have played an important role in advocating for worker interests in domestic politics. Until 2017, unions were required by law to participate in negotiations to formalize dismissals in the context of mass layoffs. [^32] In 2017, the Brazilian labor reform (Law 13,467/17) established that it was no longer mandatory for unions to take part in collective bargaining agreements during mass layoffs. We define union-affiliated parties following the classification of Ogeda et al. (2024) and compare them to all other left-wing or center parties that have no historical ties with labor unions in Brazil. [^33] Union-affiliated parties are PT, PDT, PSB, PCB, PSD, and MDB, while non-union left and center parties are CIDADANIA, PCDOB, PV, PMN, PSOL, PSTU, PCO, SD, PROS, PPL, PMB, REDE, UP, PSDB, PTB, and AVANTE. Our results show an increase in the probability of individuals becoming members or local candidates in both union-affiliated and non-union-affiliated parties (columns (7) and (8) of Table 2). In quantitative terms, we estimate a 7.5% (11.3%) increase in the probability of newly affiliating with union-aligned (non-union aligned) parties and a 15.4% (16.7%) increase in candidacy probabilities in union-affiliated (non-union affiliated) parties following mass layoffs. Thus, we do not find evidence for a large role of unions in politically mobilizing laid-off individuals.
@@ 6 Effects of Unemployment Insurance on Political Entry
The results presented in the previous section show that job loss significantly increases individuals’ likelihood of joining a party and running for local office. The heterogeneity analyses suggest that the rise in political engagement can – at least partly – be explained by the implied reduction in private-sector opportunity costs following a layoff. The opportunity cost for displaced individuals may also vary depending on their eligibility for unemployment insurance benefits. In Brazil, workers who are dismissed without just cause can receive UI benefits for 3 to 5 months, with an average replacement rate of about 80%. [^34] The length of UI payments depends on the length of employment in the 36 months before the layoff, and the replacement rate depends on the previous earnings (varying between 100% for individuals earning the minimum wage and 67% at the benefit cap, which is at 2.65 times the minimum wage). Benefit receipt mitigates the income losses from layoffs and may thereby reduce the need to seek salary from a political position. At the same time, during benefit receipt dismissed workers have lower incentives to find reemployment. [^35] The negative effects of UI eligibility on formal employment have been widely documented, including for Brazil (Gerard et al., 2020; Gerard and Gonzaga, 2021; Britto et al., 2022b). Britto (2022) also shows that an extension of the UI eligibility period in Brazil decreased total employment when taking into account the positive effect on informal employment. The negative employment effects of UI eligibility decrease labor earnings, which may offset the benefit payments, and increase individuals’ time availability. The increased time resources, in turn, may enable individuals to engage with politics. In this section, we thus investigate if eligibility for UI benefits affects laid-off individuals’ decision to enter politics. This analysis can help shed light on the mechanisms driving the effect of job loss on political entry.
@@ 6.1 Empirical strategy
As in the previous section, the estimation sample consists of all workers aged 25 to 50 displaced from a full-time, open-ended, private-sector job of at least six months tenure. To increase statistical power, we extend our layoff period to cover the years 2006 to 2014.383838We start in 2006 because this allows us to observe past layoffs up to 24 months prior to the current layoff (our employment data begin in 2004) and to measure placebo political outcomes in the two cycles prior to the current layoff (our political data begin in the 2000 cycle). We stop in 2014 because of the numerous changes introduced to the UI system after that year. This allows us to measure political outcomes over two election cycles after the layoff and labor market outcomes over up to 4 years after the layoff. [^39] While we must restrict the sample to individuals with unique names for the party membership outcome, we do not impose this restriction when analyzing candidacies, which are less common and thus require larger sample sizes. In addition, we exclude from the sample all workers who were displaced on the first or the last day of each month in their last layoff. Appendix Figure C.1 shows that the number of dismissals is higher at the very beginning and the very end of the month. Dropping these observations ensures that our RD cutoff does not coincide with spikes in the density of dismissals that occur every 30 days, i.e., that are not specific to the 16-month cutoff.
@@ 6.2 Results
Given that earlier we have documented differential effects of job loss on political entry by individuals’ education, Panels B and C of Table 3 also show the effects of UI benefits separately for low-educated workers ($<$ 12 years of schooling) and high-educated workers ($\geq$ 12 years of schooling). We find positive effects on party memberships and candidacies among the highly educated but no significant impact among the low educated. The effects on workers with high levels of education are statistically significant and meaningful in size, indicating a 6.2% increase in the likelihood of becoming a party member and a 13.0% increase in the likelihood of running for office (both relative to the control mean). These patterns are consistent with the larger effects of layoffs on individuals with higher levels of education that we presented earlier. Eligibility for UI benefits reduces individuals’ reemployment incentives, thereby increasing the time resources that they have available to engage in politics. As argued before, this shock to individuals’ opportunity costs is likely more relevant for highly educated workers, which may explain the larger effects of both job loss and UI eligibility on their decision to enter politics.
@@ 7 Implications for Political Selection
Our main results show that job loss increases individuals’ likelihood of becoming a party member or a local political candidate. An important question that follows is how these new entrants affect the political system. The increased number of party members and candidates means that parties have a larger pool from which they can select competent politicians for leadership roles. In addition, layoffs may drive individuals with different motivations and different skill levels into politics, and thereby change the average quality of politicians. Our theoretical framework in Section 2.3 has shown that layoffs can have positive consequences for political selection if private-sector wages and expected benefits from political involvement are increasing in individuals’ ability level. In particular, if layoffs are a larger shock to the opportunity costs of more competent individuals, they may improve the average competence of the candidate pool.
@@ 7.1 Individual-level evidence
To shed light on the implications of layoffs for the selection of political entrants, we start by comparing characteristics of treated and control individuals who become party members or candidates after 2012, i.e., after our layoff period. The idea of this simple comparison is as follows. Due to our matching procedure, in the full sample, treated and control individuals have identical pre-layoff characteristics (see Table 1). Any difference that we see in the pre-layoff characteristics of treated and control individuals who start political life after the layoff period should then only be driven by the effects of job loss varying across individuals. The comparison can thus tell us how individuals who enter politics due to a layoff are different from all other individuals entering politics in our sample.
We study an individual’s qualification for a political career using three different measures. First, we consider years of education as a standard measure of competence. [^42] For Brazil, Melo and Tigre (2024) provide evidence that more educated mayors are more effective managers of public resources, as they commit fewer municipal administration infringements. At the same time, they are not found to engage in fewer corrupt practices or improve the provision of local public goods. Second, we examine individuals’ monthly earnings before the layoff, which, following Dal Bó et al. (2013), can be treated as a measure of skill if the private sector rewards highly skilled workers with higher earnings. Third, we follow Besley et al. (2017) and Dal Bó et al. (2017) and consider residuals from a Mincer earnings regression. Specifically, we regress log monthly earnings on dummies for gender, education categories, age deciles, tenure deciles, firm size deciles, and municipality $\times$ two-digit industry categories. The standardized residual from this regression measures the earnings premium that a worker receives relative to observably similar workers and thus likely reflects the worker’s ability. [^43] Besley et al. (2017) show, for Sweden, that this measure positively predicts individuals’ cognitive and leadership ability, as well as electoral and policy success.
Table 4: Characteristics of new party members and candidates
(1) (2) (3) (4) (5) (6) (7) (8)
New party members Candidates
Treated Control Std Diff P-Val Treated Control Std Diff P-Val
Years of education 10.79 10.66 0.05 0.00 10.79 10.67 0.05 0.01
Earnings 2169.33 1981.66 0.10 0.00 2145.66 1956.10 0.11 0.00
Earnings residual 0.03 -0.02 0.05 0.00 0.04 0.00 0.04 0.03
Previous candidacy (%) 3.23 2.53 0.04 0.00 12.93 11.09 0.06 0.00
Observations 46,828 39,901 7,914 6,477
  • Note: The table reports the average pre-layoff characteristics of treated and control workers who newly register as a party member and run for local councilor in any post-layoff cycle (2012, 2016, or 2020). “Earnings” are monthly formal labor earnings in BRL (CPI 2018) before the layoff. “Earnings residual” is the standardized residual from a regression of log monthly earnings on dummies for gender, education categories, age deciles, tenure deciles, firm size deciles, and municipality $\times$ two-digit industry categories. “Previous candidacy” is a dummy for whether an individual ran for municipal councilor in any pre-layoff cycle (2000, 2004, or 2008). Columns (3) and (6) report standardized differences in characteristics between treated and control workers. Columns (4) and (8) report p-values for the t-tests for equality of means.
Table 4 shows the results. We observe that among individuals who become a party member or councilor candidate after the layoff, those from the treatment group exhibit higher qualifications than those from the control group. This is consistently found across all three competence measures. For example, individuals in the treatment group who run for office have earned about 9.5% more than control individuals running for office. We also consider an individual’s political experience, which we measure by whether the individual has already run for councilor in any previous election cycle (2000, 2004, or 2008). We find that among those running for office after the layoff, the share of individuals who have already run for office before the layoff is 1.9 percentage points (17%) higher in the treated compared to the control group. In sum, these results suggest that job loss increases political entry more strongly among more competent and politically experienced individuals.

### v3

@@ 4.5 Party disaggregations
Job loss may not only affect overall political participation but also lead to changes in individuals’ support of different political parties. We, therefore, study heterogeneities in the effects on memberships and candidacies by party characteristics, focusing on parties’ ideology, local incumbency, and alignment with labor unions. Estimates from the static DiD model (2) are presented in Table 2, while dynamic event-study results from model (1) are shown in Appendix Figures B.6 - B.8. We also report results from model (2) separately for each of Brazil’s major parties in Appendix Figure B.9.
@@
Ideology. Layoffs may lead individuals to align with parties whose political programs better address workers’ revised economic interests, for example, through redistributive policies. Moreover, even in the absence of economic platforms, extreme parties may be more successful in exploiting individuals’ resentment of the establishment following economic downturns (e.g., Algan et al., 2017; Autor et al., 2020). To analyze whether layoffs lead to ideological shifts, we distinguish left, center, and right parties, following the classification of Colonnelli et al. (2022). [^32] The ideological classification of parties is shown in Appendix Table B.5. Our results, shown in columns (2) to (4) of Table 2, indicate that job loss leads to an increase in party memberships and candidacies across the ideological spectrum. Relative to the control mean, we find an increase of 8.4%, 11.0% and 17.6% in the likelihood of individuals newly affiliating with left, center and right parties, respectively. The effects on candidacy reveal a similar pattern, with likelihood increases of 15.1%, 17.3%, and 21.3% of running for office in a left, center, or right-wing party. Although we observe larger effects for right-wing parties, our results show that center and left-wing parties also receive more members and candidates following layoffs. Interestingly, the event-study results presented in Appendix Figure B.6 show that the higher increase in affiliations with right-wing parties is driven only by the 2020 cycle, the first local election after Bolsonaro took power in 2019. The rise in right-wing support following economic hardships, which is well documented in the literature (e.g., Colantone and Stanig, 2018; Dippel et al., 2022; Barros and Santos Silva, 2025), appears to be contingent on the current political climate.
@@
Local incumbency. Next, we examine whether individuals punish the party of the local incumbent government after being laid off. Retrospective voting has been widely documented to rationalize the behavior of voters following economic shocks. In particular, individuals may blame incumbent politicians for local economic policies that result in mass layoffs, or for insufficient support policies that mitigate economic hardship after a layoff (e.g., Tilley et al., 2018; Ahlquist et al., 2020). In columns (5) and (6) of Table 2, we compare the effects for the party of the elected mayor in each individual’s municipality before the layoff in 2008 and all other non-incumbent parties. The results show that job loss leads to increased participation in both incumbent and non-incumbent parties. Relative to the control mean, the effects are considerably stronger for incumbent parties, with an increase of 24.0% in the probability of being affiliated with the party of the incumbent and an increase of 43.7% in the likelihood of running for office in the party of the incumbent (as opposed to 11.6% and 16.7% in non-incumbent parties). The results do not support incumbent punishment as a main driver of our results. Instead, running under the incumbent party may raise the likelihood of electoral success, thereby increasing the expected benefits of participation (consistent with our theoretical framework in Section 2.3).
@@
Union affiliation. Lastly, we investigate whether the effects depend on parties’ alignment with labor unions. Unions can influence employees’ political behavior through various activities, such as educating, informing, and mobilizing their members, endorsing selected candidates and policies, and organizing get-out-the-vote campaigns (e.g., Kim and Margalit, 2017; Matzat and Schmeißer, 2023). In Brazil, labor unions have historically been linked to left and centrist parties and have played an important role in advocating for worker interests in domestic politics. Until 2017, unions were required by law to participate in negotiations to formalize dismissals in the context of mass layoffs. [^33] In 2017, the Brazilian labor reform (Law 13,467/17) established that it was no longer mandatory for unions to take part in collective bargaining agreements during mass layoffs. We define union-affiliated parties following the classification of Ogeda et al. (2025) and compare them to all other left-wing or center parties that have no historical ties with labor unions in Brazil. [^34] Union-affiliated parties are PT, PDT, PSB, PCB, PSD, and MDB, while non-union left and center parties are CIDADANIA, PCDOB, PV, PMN, PSOL, PSTU, PCO, SD, PROS, PPL, PMB, REDE, UP, PSDB, PTB, and AVANTE. Our results show an increase in the probability of individuals becoming members or local candidates in both union-affiliated and non-union-affiliated parties (columns (7) and (8) of Table 2). In quantitative terms, we estimate a 7.5% (11.3%) increase in the probability of newly affiliating with union-aligned (non-union-aligned) parties and a 15.4% (16.7%) increase in candidacy probabilities in union-affiliated (non-union-affiliated) parties following mass layoffs. Thus, we find no evidence that unions play a major role in politically mobilizing laid-off individuals.
@@ 5 Effects of Unemployment Insurance on Political Entry
The main results presented in the previous section show that job loss significantly increases individuals’ likelihood of joining a party and running for local office. As discussed in Section 2.3, these effects may be explained by two mechanisms: income losses that reduce the opportunity costs of political entry, and increased time resources that allow individuals to engage more intensively in politics. Both channels may be influenced by individuals’ eligibility for unemployment insurance benefits. In Brazil, workers who are dismissed without just cause can receive UI benefits for 3 to 5 months, with an average replacement rate of about 80%. [^35] The length of UI payments depends on the length of employment in the 36 months before the layoff, and the replacement rate depends on the previous earnings (varying between 100% for individuals earning the minimum wage and 67% at the benefit cap, which is at 2.65 times the minimum wage). Benefit receipt mitigates the income losses from layoffs and may thereby reduce the financial incentives to pursue a political position. At the same time, during benefit receipt dismissed workers have lower incentives to find reemployment. [^36] The negative effects of UI eligibility on formal employment have been widely documented, including for Brazil (Gerard et al., 2020; Gerard and Gonzaga, 2021; Britto et al., 2022b). Britto (2022) also shows that an extension of the UI eligibility period in Brazil decreased total employment when taking into account the positive effect on informal employment. The negative employment effects of UI eligibility decrease labor earnings, which may offset the benefit payments, and increase individuals’ time availability. The increased time resources, in turn, may facilitate individuals’ political engagement. In this section, we investigate whether eligibility for UI benefits affects laid-off individuals’ decision to enter politics, which can help shed light on the mechanisms driving the effect of job loss on political entry.
@@ 5.1 Empirical strategy
As in the previous section, the estimation sample consists of all workers aged 25 to 50 displaced from a full-time, open-ended, private-sector job of at least six months tenure. To increase statistical power, we extend our layoff period to cover the years 2006 to 2014.393939We start in 2006 because this allows us to observe past layoffs up to 24 months prior to the current layoff (our employment data begin in 2004) and to measure placebo political outcomes in the two cycles prior to the current layoff (our political data begin in the 2000 cycle). We stop in 2014 because of the numerous changes introduced to the UI system after that year. This allows us to measure political outcomes over two election cycles after the layoff and labor market outcomes over up to 4 years after the layoff. [^40] While we must restrict the sample to individuals with unique names for the party membership outcome, we do not impose this restriction when analyzing candidacies, which are less common and thus require larger sample sizes. In addition, we exclude from the sample all workers who were displaced on the first or the last day of each month in their last layoff. Appendix Figure C.1 shows that the number of dismissals is higher at the very beginning and the very end of the month. Dropping these observations ensures that our RD cutoff does not coincide with spikes in the density of dismissals that occur every 30 days, i.e., that are not specific to the 16-month cutoff.
@@ 5.2 Results
Given that earlier we have documented differential effects of job loss by education, Panels B and C of Table 3 also show the effects of UI benefits separately for low-educated workers ($<$ 12 years of schooling) and high-educated workers ($\geq$ 12 years of schooling). We find positive effects on party memberships and candidacies among the highly educated but no significant impact among the low educated. The effects on workers with high levels of education are statistically significant and meaningful in size, indicating a 6.2% increase in the likelihood of becoming a party member and a 13.0% increase in the likelihood of running for office (both relative to the control mean). These findings are consistent with the time-resources mechanism: by reducing reemployment incentives, eligibility for UI benefits frees up time that can be redirected toward political engagement. As suggested by our theoretical framework in Section 2.3, the importance of this channel increases with the expected benefits of political participation. These benefits may vary with education – for example, if more educated individuals have greater chances of electoral success or are more likely to leverage political connections for gaining access to public-sector jobs. Such mechanisms can explain why both job loss and UI eligibility have stronger effects on political entry among the highly educated.
@@ 6 Implications for Political Selection
Our main results show that job loss increases individuals’ likelihood of becoming a party member or a local political candidate. An important question that follows is how these new entrants affect the political system. The increased number of party members and candidates means that parties have a larger pool from which they can select competent politicians for leadership roles. In addition, layoffs may drive individuals with different motivations and different skill levels into politics, and thereby change the average quality of politicians. As discussed in Section 2.3, layoffs can have positive consequences for political selection if income losses or the expected returns from political entry increase with individuals’ ability level. In particular, if more competent individuals experience greater reductions in private-sector opportunity costs or anticipate higher benefits from political engagement during unemployment, layoffs may raise the average competence of the candidate pool.
@@ 6.1 Individual-level evidence
To explore the impact of layoffs on political selection, we examine the individual characteristics of candidates and political members who enter politics as a result of being laid off. We use a combination of sample moments in an instrumental variable framework to estimate the average characteristics of three groups: (1) compliers—individuals who enter politics because of the layoff, (2) always-takers—those who enter politics regardless of whether they are laid off or not, and (3) never-takers—those who do not enter politics regardless of their layoff status. To identify compliers, we build on the kappa weighting approach introduced by Abadie (2003). We estimate the expected value of covariate $X_{i}$ for compliers as
$E[X_{i}|compliers]=\frac{E[\kappa_{i}X_{i}]}{E[\kappa_{i}]},$
where $\kappa_{i}$ is defined as
$\kappa_{i}=Y_{i}\frac{Treat_{i}-P(X_{i})}{P(X_{i})(1-P(X_{i}))}.$
Here, $Y_{i}$ indicates whether an individual enters politics in a post-layoff cycle (either as a party member or a councilor candidate), and $Treat_{i}$ represents their layoff status. We estimate the propensity score, $P(X_{i})=P(Treat_{i}=1|X_{i})$, using the set of all covariates considered in the complier analysis. [^43] Słoczyński et al. (2025) provide a comprehensive treatment of various kappa weighting estimators that are implemented in the Stata command kappalate. We use the estimator based on their $\kappa_{i1}$ definition. The authors recommend these kappa weights for applications without always-takers (one-sided noncompliance) to avoid near-zero denominators. Similarly, in our case, the number of always-takers is very small relative to the number of never-takers.
The complier analysis relies on two main assumptions. First, we assume instrument independence. This is supported by the use of the mass-layoff sample described in Section 4, in which treated and control workers are matched on various pre-layoff characteristics (see, again, Table 1 for balance checks). Moreover, we account for any remaining differences in covariates in the propensity score estimation. Second, we assume treatment monotonicity. In our context, this means that individuals are not discouraged from entering politics due to a layoff. While we cannot entirely rule out this possibility, the heterogeneity analyses across individual, local, and party characteristics reveal significant positive effects across all considered subgroups, suggesting that the presence of defiers – if any – is likely limited (see, again, Figure 5, Appendix Figure B.4, and Table 2).
When analyzing the characteristics of political entrants, we focus on measures that capture individual competence, in addition to socio-demographic characteristics. First, we consider years of education as a standard measure of competence. [^44] For Brazil, Melo and Tigre (2024) provide evidence that more educated mayors are more effective managers of public resources, as they commit fewer municipal administration infringements. However, they are not found to engage in fewer corrupt practices or improve the provision of local public goods. Second, we examine individuals’ log hourly wage before the layoff, which, following Dal Bó et al. (2013), can be treated as a measure of skill, as the private sector typically rewards highly skilled workers with higher wages. Third, we consider person wage effects from an AKM model that regresses log wages on person fixed effects and firm fixed effects (Abowd et al., 1999). The person fixed effects measure the wage premia that workers receive independent of the firm that they work at, likely reflecting workers’ ability that is rewarded across all firms. [^45] To our knowledge, we are the first to use AKM effects to study the competence of politicians. The AKM model is fitted on all RAIS contracts in the pre-layoff years 2004 to 2008. Appendix Table D.1 reports a summary of the AKM estimation results. Moreover, we consider residualized AKM person effects, controlling for gender, race, age deciles, education categories, as well as the full set of interactions between these variables. The resulting residuals capture wage premia relative to observably similar workers. [^46] Besley et al. (2017) and Dal Bó et al. (2017) also consider residuals from Mincer wage regressions. Our AKM measures additionally purge wages from firm variation in wage-setting, which may not reflect individual ability. To validate our competence measures, we correlate them with local councilors’ misconduct detected during the electoral process and with their electoral performance. Results are reported in Appendix Table D.2 and show that all our competence measures are significantly negatively correlated with the probability of misbehaving in the electoral process and significantly positively related the the probability of being elected. [^47] The misbehavior variable captures a broad range of irregularities detected by the Brazilian Electoral Court during the elections in 2016 and 2020 that resulted in candidacy annulment. These include abuse of power, failure to meet registration requirements, vote buying, prohibited conduct, ineligibility under the “Clean Record Act” (Lei da Ficha Limpa), illicit use of campaign resources, objections to the candidacy, rejection by the party or coalition, and invalidation of the party. Finally, in addition to the competence measures, we also study socio-demographic representation in terms of gender and age.
Figure 9 presents the results. Overall, we find that individuals who enter politics because of a layoff (compliers) tend to be positively selected in terms of competence relative to always-takers and never-takers. For example, layoff-induced councilor candidates have an average of 12.0 years of education, compared to about 10.9 years among individuals who always or never run for local
Figure 9: Characteristics of political entrants
A. New party membership
B. Candidacy
Note: The figure shows the average pre-layoff characteristics of individuals in our estimation sample who are compliers (those induced to enter politics due to the layoff), always-takers (those who enter politics regardless of being laid off or not), and never-takers (those who do not enter politics regardless of being laid off or not). Panel A and B report results separately for our two outcomes: registering as a party member and running for local councilor in a post-layoff cycle (2012, 2016, and 2020). While the estimation sample is restricted to individuals with unique names for the party membership outcome, we do not impose this restriction when analyzing candidacies. For the characteristics provided by the TSE, we also show the average characteristics of the total population of candidates (also in 2012, 2016, and 2020). In Panel A, $N$ = 11,389 compliers, 70,088 always-takers, and 4,868,859 never-takers. In Panel B, $N$ = 3,350 compliers, 24,227 always-takers, 9,855,945 never-takers, and 1,428,485 candidates in the total population. Log wages and AKM person effects are measured based on individuals’ contracts in RAIS between 2004 and 2008, and we also show residualized AKM person effects that control for gender, race, age deciles, education categories, as well as the full set of interactions between these variables. AKM person effects and residuals are standardized across all workers in RAIS between 2004 and 2008.
office, regardless of whether they are laid off or not. [^48] Interestingly, we observe relatively small differences between always-takers and never-takers on all measures of competence, suggesting that political selection may be weak, on average. For layoff-induced party members, we find similarly higher education, compared to always- and never-takers. Moreover, there are clear differences in log wages and AKM person effects (and their residuals) that also indicate higher competence among compliers. For instance, layoff-induced candidates have AKM person effects about 0.2 standard deviations higher than those of always- and never-takers, and for layoff-induced party members the difference is even larger.
These results suggest that layoffs improve the competence of political entrants relative to always-takers in our estimation sample. To better assess how layoff-induced entrants can change the overall pool of politicians – beyond our estimation sample – we can compare our complier estimates to information on all councilor candidates reported by the TSE. The average education of all councilor candidates between 2012 and 2020 is 11.1 years, showing that layoff-induced candidates have about one more year of education than the average candidate. [^49] Unfortunately, for the full candidate population, we observe only education and other socio-demographic characteristics. Wage data are unavailable in the TSE data, and similar information is not provided for the population of party members.
Importantly, the higher levels of competence do not come at the expense of reduced representation. Our results indicate that individuals who run for office because of the layoff are more likely to be female and younger than always-takers and the full candidate population. [^50] For party members, compliers are also more likely to be female, but slightly older than always-takers. In sum, these findings suggest that job loss increases political entry more strongly among individuals with higher competence, and is also accompanied by improved political representation.